# Properties of current sheets in two-dimensional tearing-mediated magnetohydrodynamic turbulence

### v1

@@ 3.3 Case study of a current sheet
In summary, we conclude that this single current sheet undergoes the following stages: (1) Formation by thinning and lengthening accompanied by laminar reconnection. (2) Growth of linear tearing instability when the aspect ratio approaches the Sweet-Parker scaling $a/L\sim S_{L}^{-1/2}$. (3) Fast dissipation of the current sheet by nonlinear tearing instability. We note that, the Lundquist number $S_{L}=6.5\times 10^{4}$ of this current sheet is not significantly larger than the critical Lundquist number, approximately $O(10^{4})$, for the Sweet-Parker type current sheet to be tearing-unstable (e.g. D. Biskamp, 1986; N. Loureiro et al., 2007; C. Shi et al., 2018). In the nearly-collisionless space plasma with much larger Lundquist numbers, the growth of linear tearing mode may happen earlier, when the aspect ratio $a/L$ approaches $S_{L}^{-1/3}$ instead of $S_{L}^{-1/3}$ (F. Pucci & M. Velli, 2013; A. Tenerani et al., 2015, 2016).
3.4 Statistics of current sheets
@@
We apply the analysis described in Section 3.3 to all the identified current sheets and present the results in Figure 8. Top to bottom rows show time moments $t=0.15$, $0.3$, and $0.6$ respectively. Left column shows aspect ratio $a/L$ versus Lundquist number $S_{L}$, middle column shows upstream magnetic field $B_{up}$ versus current sheet thickness $a$, and right column shows current sheet length $L$ versus thickness $a$. Each dot corresponds to one identified current sheet, and the dot is color-coded by the peak current density in the current sheet. In the left column, we plot reference lines $a/L\propto S_{L}^{-1/2}$, i.e. the Sweet-Parker type current sheet model. We can see that for all the three moments, the data points show a negative correlation between $a/L$ and $S_{L}$, and they roughly obey $a/L\sim S_{L}^{-1/2}$. In the middle and right columns, we plot reference lines $B_{up}\propto a^{1/4}$ and $L\propto a^{3/4}$, which are predictions by the SDDA theory (S. Boldyrev, 2006). There is a strong positive correlation between $L$ and $a$, and they may roughly follow the power-law relation $L\propto a^{3/4}$. In contrast, there is no clear correlation between $B_{up}$ and $a$. We note that, at $t=0.15$, i.e. before tearing instability is triggered, most of the current sheets have lengths close to the wavelengths of initial fluctuations marked by the yellow shades. Moreover, panel (b1) shows that the upstream magnetic field is also quite consistent among all the current sheets, slightly above the initial fluctuation level ($\delta b\approx 0.17$). These results indicate that the generation of current sheets is dominated by the largest turbulence eddies in the system.
@@
4 Discussion: Current sheets and scale-dependent dynamic alignment
@@
and similar for $\theta_{z^{\pm}}$. We employ a Monte-Carlo method to evaluate the angles. That is, we randomly select two grid points $\bm{x_{1}}$ and $\bm{x_{2}}$ such that $\bm{l}=\bm{x_{2}}-\bm{x_{1}}$ and define $\delta\bm{b}=\bm{b}(\bm{x_{2}})-\bm{b}(\bm{x_{1}})$ and same for $\delta\bm{u}$ and $\delta\bm{z^{\pm}}$. We repeat this random process for $10^{8}$ times and calculate quantities like $\langle\left|\delta\bm{u}\times\delta\bm{b}\right|\rangle$, $\langle|\delta\bm{u}||\delta\bm{b}|\rangle$, and subsequently the alignment angles for the binned $l$ and $\phi$. Figure 10 confirm that both the alignment angles are smaller as we go toward smaller scales. In the classic SDDA models (S. Boldyrev, 2006; N. F. Loureiro & S. Boldyrev, 2017), the alignment angle scales as $\theta_{ub}\propto\lambda^{0.25}$, while in the later model by B. D. Chandran et al. (2015) the angles follow $\theta_{ub}\propto\lambda^{0.21}$ and $\theta_{z^{\pm}}\propto\lambda^{0.1}$. Although these previous results are not directly applicable to the 2D turbulence, we still plot two reference lines with slopes of 0.25 and 0.10 respectively in each panel of Figure 10 for visual assistance. Panels (a) and (b) show that, at early stage $t=0.15$, a single power law is observed for both the angles, and the slope for $\theta_{z^{\pm}}$ is steeper than that for $\theta_{ub}$. At later times, double power-law relations form for both angles with a break scale at $l\sim(1-2)\times 10^{-2}$. The large scale shows a slope close to $0.1$ and the small scale shows a slope close to $0.25$. Panels (c) and (d) also show a clear break at $\lambda\sim(1-2)\times 10^{-2}$. The power-law slopes for small scales are similar to the top panels, but the slopes for large scales are quite different from the top panels. What controls this break frequency and how the alignment angles change with scale is still unclear and requires future investigation. Another noteworthy point is that, for $\theta_{z^{\pm}}(l)$, the alignment angle first evolves to a low level at $t=0.15$ and then gradually rises. $\theta_{z^{\pm}}(\lambda)$ shows a similar evolution trend, though only for small scales. This is reminiscent of the evolution of Kurtosis($\bm{b}$) shown in Figure 4 and the residual energy shown in Figure 1 and it implies that $\theta_{z^{\pm}}$ is likely related to the evolution of turbulence current sheets.
@@
Last, we quantify the anisotropy of the turbulence “eddies” using the method described in (N. Sioulas et al., 2024b). We apply the Monte-Carlo method to evaluate the second-order structure function of magnetic field while taking $\phi$ into consideration. As shown by Figure 11(a), $S_{2}(\bm{b})$ is clearly anisotropic between the directions parallel ($\xi$, $\phi=0$) and perpendicular ($\lambda$, $\phi=90^{\circ}$) to the local magnetic field increment. We numerically determine the function $\xi(\lambda)$ by equating $S_{2}(\bm{b},\xi)$ and $S_{2}(\bm{b},\lambda)$, and the result is shown in panel (b), which displays $\xi$ versus $\lambda$, and panel (c), which displays $\xi/\lambda$ versus $\lambda$. In general, $\xi>\lambda$, even at $t=0$, meaning that the turbulence eddies are stretched along the polarization direction. $\lambda/\xi$ is roughly scale-independent at $t=0$, with values 0.5-0.6, but becomes scale-dependent soon after the simulation starts. Similar to the Kurtosis shown in Figure 4 and $\theta_{z^{\pm}}$ shown in Figure 10, $\lambda/\xi$ drops to its lowest value at $t=0.15$ and then starts to rise, implying that it is related to the formation of turbulence current sheets. However, we emphasize that, a temporal correlation or concurrence does not ensure causal relation. The fact that current sheets, dynamic alignment, and turbulence eddy anisotropy evolve in a similar pattern does not mean that the current sheets are generated as a result of the dynamic alignment and eddy anisotropy. In fact, the estimated $\lambda/\xi$ (Figure 11) are much larger than the aspect ratio of the current sheets (Figure 8(a1-a3)). In addition, the filling factor of the current sheets is very low (<0.1<0.1, Figure 5) while the alignment angles and eddy anisotropy are evaluated over the whole simulation domain, meaning that the current sheet properties have negligible contribution in evaluation of the alignment angles and eddy anisotropy.
5 Summary
@@
    As the current sheets continue to thin, tearing instability onsets when the current sheet aspect ratio approaches $S^{-1/2}$.
@@
CS is supported by NSF SHINE #2229566 and NASA ECIP #80NSSC23K1064. The numerical simulations were conducted on Derecho: HPE Cray EX System (https://doi.org/10.5065/qx9a-pg09) of Computational and Information Systems Laboratory (CISL), National Center for Atmospheric Research (NCAR) (allocation No. UCLA0063).

### v2

@@ III.3 Case study of a current sheet
In summary, we conclude that this single current sheet undergoes the following stages: (1) Formation by thinning and lengthening accompanied by laminar reconnection. (2) Growth of linear tearing instability when the aspect ratio approaches the Sweet-Parker scaling $a/L\sim S_{L}^{-1/2}$. (3) Fast dissipation of the current sheet by nonlinear tearing instability. We note that, the Lundquist number $S_{L}=6.5\times 10^{4}$ of this current sheet is not significantly larger than the critical Lundquist number, approximately $O(10^{4})$, for the Sweet-Parker type current sheet to be tearing-unstable (e.g. biskamp1986magnetic; loureiro2007instability; shi2018marginal). In the nearly-collisionless space plasma with much larger Lundquist numbers, the growth of linear tearing mode may happen earlier, when the aspect ratio $a/L$ approaches $S_{L}^{-1/3}$ instead of $S_{L}^{-1/2}$ (pucci2013reconnection; tenerani2015magnetic; tenerani2016ideally; del2016ideal).
III.4 Statistics of current sheets
@@
We apply the analysis described in Section III.3 to all the identified current sheets and present the results in Figure 8. Top to bottom rows show time moments $t=0.15$, $0.3$, and $0.6$ respectively. Left column shows aspect ratio $a/L$ versus Lundquist number $S_{L}$, middle column shows upstream magnetic field $B_{up}$ versus current sheet thickness $a$, and right column shows current sheet length $L$ versus thickness $a$. Each dot corresponds to one identified current sheet, and the dot is color-coded by the peak current density in the current sheet. In the left column, we plot reference lines $a/L\propto S_{L}^{-1/2}$, i.e. the Sweet-Parker type current sheet model. We can see that for all the three moments, the data points show a negative correlation between $a/L$ and $S_{L}$, and they roughly obey $a/L\sim S_{L}^{-1/2}$. In the middle column, we plot reference lines $B_{up}\propto a^{1/4}$, which is the prediction by the SDDA theory (boldyrev2006spectrum). One can see that there is no clear correlation between $B_{up}$ and $a$. In contrast, the right column shows a strong positive correlation between $L$ and $a$. For reference, we plot $L\propto a^{2}$, which is the scaling relation given by Sweet-Parker model, i.e. $a/L\sim(LB_{up}/\eta)^{-1/2}$, assuming that the upstream Alfvén speed is independent of $a$. Similarly, for ideal-tearing scaling relation $a/L\sim(LB_{up}/\eta)^{-1/3}$, we will get $L\propto a^{3/2}$. Note that, SDDA theory predicts $L\propto a^{3/4}$, much shallower than what is shown here. We point out that, at $t=0.15$, i.e. before tearing instability is triggered, most of the current sheets have lengths close to the wavelengths of initial fluctuations marked by the yellow shades. Moreover, panel (b1) shows that the upstream magnetic field is also quite consistent among all the current sheets, slightly above the initial fluctuation level ($\delta b\approx 0.17$). These results indicate that the generation of current sheets is dominated by the largest turbulence eddies in the system (papini2019can; khan2025does), and their subsequent evolution is dominated by the recursive tearing instability (shibata2001plasmoid; tenerani2015magnetic).
@@
IV Discussion: Current sheets and scale-dependent dynamic alignment
@@
and similar for $\theta_{z^{\pm}}$. We employ a Monte-Carlo method to evaluate the angles. That is, we randomly select two grid points $\bm{x_{1}}$ and $\bm{x_{2}}$ such that $\bm{l}=\bm{x_{2}}-\bm{x_{1}}$ and define $\delta\bm{b}=\bm{b}(\bm{x_{2}})-\bm{b}(\bm{x_{1}})$ and same for $\delta\bm{u}$ and $\delta\bm{z^{\pm}}$. We repeat this random process for $10^{8}$ times and calculate quantities like $\langle\left|\delta\bm{u}\times\delta\bm{b}\right|\rangle$, $\langle|\delta\bm{u}||\delta\bm{b}|\rangle$, and subsequently the alignment angles for the binned $l$ and $\phi$. Figure 10 confirm that both the alignment angles are smaller as we go toward smaller scales. In the classic SDDA models (boldyrev2006spectrum; loureiro2017role), the alignment angle scales as $\theta_{ub}\propto\lambda^{0.25}$, while in the later model by chandran2015intermittency the angles follow $\theta_{ub}\propto\lambda^{0.21}$ and $\theta_{z^{\pm}}\propto\lambda^{0.1}$. Although these previous results are not directly applicable to the 2D turbulence, we still plot two reference lines with slopes of 0.25 and 0.10 respectively in each panel of Figure 10 for visual assistance. Panels (a) and (b) show that, at early stage $t=0.15$, a single power law is observed for both the angles, and the slope for $\theta_{z^{\pm}}$ is steeper than that for $\theta_{ub}$. At later times, double power-law relations form for both angles with a break scale at $l\sim(1-2)\times 10^{-2}$. The large scale shows a slope close to $0.1$ and the small scale shows a slope close to $0.25$. Panels (c) and (d) also show a clear break at $\lambda\sim(1-2)\times 10^{-2}$. The power-law slopes for small scales are similar to the top panels, but the slopes for large scales are quite different from the top panels. What controls this break frequency and how the alignment angles change with scale is still unclear and requires future investigation. Another noteworthy point is that, for $\theta_{z^{\pm}}(l)$, the alignment angle first evolves to a low level at $t=0.15$ and then gradually rises. $\theta_{z^{\pm}}(\lambda)$ shows a similar evolution trend, though only for small scales. This is reminiscent of the evolution of Kurtosis($\bm{b}$) shown in Figure 4 and the residual energy shown in Figure 1 and it implies that $\theta_{z^{\pm}}$ is likely affected by the evolution of turbulence current sheets.
@@
Last, we quantify the anisotropy of the turbulence “eddies” using the method described in (sioulas2024higher). We apply the Monte-Carlo method to evaluate the second-order structure function of magnetic field while taking $\phi$ into consideration. As shown by Figure 11(a), $S_{2}(\bm{b})$ is clearly anisotropic between the directions parallel ($\xi$, $\phi=0$) and perpendicular ($\lambda$, $\phi=90^{\circ}$) to the local magnetic field increment. We numerically determine the function $\xi(\lambda)$ by equating $S_{2}(\bm{b},\xi)$ and $S_{2}(\bm{b},\lambda)$, and the result is shown in panel (b), which displays $\xi$ versus $\lambda$, and panel (c), which displays $\xi/\lambda$ versus $\lambda$. In general, $\xi>\lambda$, even at $t=0$, meaning that the turbulence eddies are stretched along the polarization direction. $\lambda/\xi$ is roughly scale-independent at $t=0$, with values 0.5-0.6, but becomes scale-dependent soon after the simulation starts. Similar to the Kurtosis shown in Figure 4 and $\theta_{z^{\pm}}$ shown in Figure 10, $\lambda/\xi$ drops to its lowest value at $t=0.15$ and then starts to rise, implying that it is more or less related to the formation of turbulence current sheets. However, we emphasize that, a temporal correlation or concurrence does not ensure causal relation. The fact that current sheets, dynamic alignment, and turbulence eddy anisotropy evolve in a similar pattern does not mean that the current sheets are generated as a result of the dynamic alignment and eddy anisotropy. In fact, the estimated $\lambda/\xi$ (Figure 11) are much larger than the aspect ratio of the current sheets (Figure 8(a1-a3)), which is mostly less than 0.1. In addition, the filling factor of the current sheets is very low (<0.1<0.1, Figure 5) while the alignment angles and eddy anisotropy are evaluated over the whole simulation domain, meaning that the current sheet properties should have negligible contribution in the evaluation of the alignment angles and eddy anisotropy.
V Summary
@@
    As the current sheets continue to thin, tearing instability onsets when the current sheet aspect ratio approaches $S^{-1/2}$. This process happens in a recursive way and generates smaller and smaller current sheets.
@@
CS is supported by NSF SHINE #2548299 and NASA ECIP #80NSSC23K1064. The numerical simulations were conducted on Derecho: HPE Cray EX System (https://doi.org/10.5065/qx9a-pg09) of Computational and Information Systems Laboratory (CISL), National Center for Atmospheric Research (NCAR) (allocation No. UCLA0063).